OpenStudy (anonymous):

solve the following system 2x+3x-z=1 3x+y+2z=12 x+2y-3z=-5 a) (3, 1, 2) b) (-3, 1 , 2) c) (3 , -1 , 2) d) (3 , 1 , -2)

OpenStudy (anonymous):

Are you sure you posted this correctly. You should double-check...

OpenStudy (anonymous):

2x+3y-z=1 3x+y+2z=12 x+2y-3z=-5

OpenStudy (anonymous):

I'm almost certain somebody will solve this for you...

OpenStudy (anonymous):

i really don't know who can help me with this ..

OpenStudy (anonymous):

The only reason why I don't want to help you is all you're interested in is the answer

OpenStudy (anonymous):

oh come on man, please ..

OpenStudy (anonymous):

Please give you the answer???

OpenStudy (anonymous):

or solve it, it's okay just please please help me, i need the answer .

OpenStudy (anonymous):

x=3, y=-1, z=2

OpenStudy (anonymous):

I'm not really a big fan of giving just the answer only...

OpenStudy (amistre64):

since they give you the points, its easier to see which one fits rather than trying to work out the answer

OpenStudy (amistre64):

it aint a or d

OpenStudy (amistre64):

looks like c fits all of them

OpenStudy (amistre64):

when given multiple choice, the answer is already there, just see what fits :)

OpenStudy (anonymous):

(2x + 3y -z = 1)2 ~ 4x + 6y -2z = 2 4x + 6y -2z = 2 3x+y+2z=12 ~ 7x + 7y = 14 (3x+y+2z=12)3 (x+2y-3z=-5)2 ~ 11x + 7y = 26 7x + 7y = 14 -(11x + 7y = 26) ~ -4x = -12 - - -> x=3 11(3) + 7y = 26 33 + 7y = 26 - - -> y=-1 2x + 3y - z = 1 2(3) + 3(-1) -z = 1 6 + (-3) -z = 1 3-z = 1 -z = 1-3 -z = -2 - - -> z=2

OpenStudy (anonymous):

ITS letter c